8) Let a=7a - 1 - 120,-2,4, =1, a, =2.Use the strong induction to show that for all n 20, a=2(3)- 4
Answer #1

Given that an=7 ani -12 an-2 ,201 ,,2 Shanu Charm an= = 2(3) -40? 100 & Step 1: let noo 20=213°) -4° -2619-1 - 2-1-1 >> 001K k akt = 14 E14(3) -7(40k . -.8 (3) * +3140k 6 (30k - 4 (4)k akti - =2.3 (3jk - 4140k K+1 Qk41 = 2(3) -14) K+1 => result is